# Beyond the Kolmogorov Barrier: A Learnable Weighted Hybrid Autoencoder for Model Order Reduction

### v4

@@ 1 Introduction
Computational fluid dynamics involves solving large dynamical systems with millions of degrees of freedom, resulting in significant computational overhead. In order to alleviate this problem, reduced-order modeling [1, 2] is widely used, which uses a smaller number of modes to provide an approximate solution at a lower computational expense. A crucial step in reduced-order modeling is the projection of high-dimensional system states to a reduced latent space [3]. The quality of projection can determine the overhead error for reduced-order modeling [4]. Linear dimensionality reduction techniques such as Proper Orthogonal Decomposition (POD) [5, 6, 7] are often used to create efficient representations of large-scale systems by projecting the solution manifold onto the space spanned by a set of linear orthonormal basis.
@@
Advances in deep learning techniques, such as deep autoencoders (AE) [8], capture intrinsic non-linear features for better compression, outperforming POD in low ranks, thus overcoming the well-known Kolmogorov barrier [9, 10]. Kolmogorov barrier is defined as a slow decay of the Kolmogorov $n$-width [11, 12], given by $d_{n}(\mathcal{M})=\inf_{\begin{subarray}{c}U_{n}\subset\mathcal{V},\dim(U_{n})=n\end{subarray}}\sup_{x(\cdot;t,\mu)\in\mathcal{M}}\inf_{\hat{x}\in U_{n}}\|x(\cdot;t,\mu)-\hat{x}\|$, where $\mathcal{M}$ is the solution manifold, $\mathcal{V}$ is the ambient Hilbert space, $U_{n}$ is any $n$-dimensional subspace of $\mathcal{V}$, $x(\cdot;t,\mu)$ is a solution to a parameterized PDE at time $t$ and parameter $\mu$, $\hat{x}$ is its projection onto $U_{n}$, and $\|\cdot\|$ denotes the norm in $\mathcal{V}$. The $n$-width quantifies the minimum worst-case projection error achievable by any $n$-dimensional linear subspace. This slow decay limits the best achievable error when using linear projection-based model reduction. Milano and Koumoutsakos [13] highlighted one of the first works to utilize a fully connected autoencoder to reconstruct the flow field, offering better performance compared to POD. Further studies have reported the usage of convolutional neural networks on 2D or 3D flow fields [14, 15, 16]. Such methods have been adopted in the fluid dynamics community to obtain a nonlinear model order reduction [17, 18], but they do not provide projection error convergence as the rank of the latent space increases. Recently, there have also been studies on using the hybrid approach [19, 20] combining POD with deep learning, by passing the latent space produced by POD to a neural network to find the corrections required to enhance reconstruction. These hybrid techniques have proven to enhance reconstruction beyond the capabilities of vanilla autoencoders. Unlike these approaches, which treat POD as a fixed preprocessing step and use neural networks only to adjust the POD output, in this work we propose a novel dimensionality reduction technique that combines traditional dimensionality reduction technique, i.e., POD, with deep learning techniques in a weighted manner at the encoder and decoder stage, where such weights of hybridization are also learned from data, to achieve a more effective dimensionality reduction. We also compare this approach with a straightforward hybrid technique, where a direct sum of POD and AE is utilized to construct the autoencoder, demonstrating the need of using learnable weighting parameters between POD and AE. Interestingly, our proposed approach obtains flat minima as opposed to other approaches, which contributes to the improved generalization and noise robustness.
@@
Building on our proposed framework, we also demonstrate the application of our framework to surrogate modeling tasks, particularly in the context of PDE system evolution. We explore two distinct applications of our framework. First, we integrate our framework with the Koopman operator, where our framework facilitates the mapping of Koopman embedding back to physical space, while learning a linear forward model to evolve the system states in a reduced space. Second, we leverage long-short-term memory (LSTM) for latent state evolution [21]. LSTMs are usually trained on the trajectories of reduced states, obtained via POD [22, 23] or a non-linear autoencoder [24, 25] to serve as a surrogate for the temporal evolution of latent states. The computational costs can be prohibitive when using LSTMs on full states, which can typically be in the range of millions for complex scenarios in fluid dynamics. By integrating our improved dimensionality reduction technique with a forward model for the latent states, we show that the overall framework leads to a more accurate prediction of the system dynamics and provide insights into the individual error contributions from the dimensionality reduction and time series model. This underscores the fact that the quality of the reduced representation is the primary bottleneck in achieving highly accurate predictions, as suboptimal dimensionality reduction inherently limit the effectiveness of any downstream models.
@@ 2.1 Dimensionality Reduction
Given the training dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{M}$, we trained the autoencoder by minimizing the mean-squared error (MSE) $\min_{\mathbf{\Theta},\mathbf{a},\mathbf{b}}\frac{1}{M}\sum_{i=1}^{M}\|\mathbf{x}_{i}-\hat{\mathbf{x}_{i}}\|^{2}$, where $\|\cdot\|$ is Frobenius norm, and $\mathbf{\Theta}$ refers to the set of the trainable parameters of neural network encoder $\phi_{\textrm{NN}}$ and decoder $\psi_{\textrm{NN}}$. $\mathbf{\Theta}$ is initialized using standard Kaiming initialization [26]. Motivated by Wang et al. [27], we choose to initialize $\mathbf{a}$ and $\mathbf{b}$ with zeros, leading to the proposed framework being equivalent to the classical POD at the beginning of neural network training. Thus, the model starts from the optimal linear encoder and becomes progressively nonlinear as the training proceeds. We choose Adam optimizer with learning rate of $10^{-4}$ for $\mathbf{\Theta}$ and $10^{-5}$ for $\mathbf{a}$ and $\mathbf{b}$.
To emphasize the role of learnable weight $\mathbf{a}$ and $\mathbf{b}$, we also implement a straightforward hybrid approach [28], which simply adds the latent states from POD and NN. Similarly, the state of the reconstructed system is the sum of the output of the POD decoder and the NN decoder. We name this hybrid approach as simple hybrid autoencoder. The difference of two approaches is further illustrated in Figure 1. As we shall see in the following sections, this comparison underscores the importance of using a weighted blend of the encoded and decoded spaces, as without it, the improvement over POD is at most incremental.
@@ 2.2 Koopman Forecasting
Koopman theory [29] states that any nonlinear dynamical system can be linearized by lifting into a (possiblly infinite-dimensional) space of observable functions. Hence, given a continuous dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$, for any observable $\xi$, one defines the Koopman operator $\mathcal{K}$ as:
@@
To overcome these challenges, Lange et al. [30] proposed Koopman forecasting,
@@ 2.3 Surrogate Modeling for Time-Dependent PDEs
Our surrogate modeling strategy involves the usage of the aforementioned techniques to obtain a latent representation and LSTM to determine its evolution over time. Training is divided into two stages. The dimensionality reduction framework is first trained to obtain a low-dimensional embedding of the system states, which subsequently serves as the training data for the time series prediction model to evolve the system in time[24, 31]. The advantage of training them separately is that it allows us to clearly demarcate the contribution of each stage to the overall surrogate model, which is particularly important as the crux of the current study is the introduction of a novel framework for dimensionality reduction.
@@ 3.1.2 Homogeneous isotropic turbulence (HIT)
Direct numerical simulation data of forced homogeneous isotropic turbulence is obtained from the Johns Hopkins turbulence database [32]. The data set is generated by solving the forced Navier-Stokes equation on a periodic cubic box using the pseudospectral method. We interpolate the velocity field, $u(x,t)=(u_{x},u_{y},u_{z})\in\mathbb{R}^{3}$, from the original dataset of resolution $1024^{3}$ to grid sizes of $16^{3}$, $32^{3}$ and $64^{3}$ and extract the data from timestep 1 to 2048 with a stride of 16, resulting around 128 snapshots.
We choose a deep convolutional autoencoder [33] as the NN part of the proposed framework. The details of the architecture pertaining to the NN based encoder and decoder and the training hyper parameters are shown in Table 1. Without shuffling, we systematically sample the dataset by selecting every alternate snapshot to curate the training and testing data, each comprising 64 snapshots.
@@ 3.2.2 Flow Over Cylinder
We utilize the time evolution data of vorticity in a two-dimensional flow over a cylinder at $Re\approx 100$[30]. The initial 50 temporal snapshots are used in training the Koopman-Decoder model, and the remaining 100 are kept out for testing. The details of the non-linear part of the decoder and training parameters are provided in Table 3. The Koopman model is parametrized by two frequencies, essentially evolving the dynamics in a latent space of rank 4.
@@ 3.3 Surrogate Modeling Datasets
A fully connected network using the same architecture as outlined in Section 3.1.1 is used for the NN based encoder and decoder. The dimensionality reduction framework was trained with a learning rate of $1\times 10{-3}$, batch size of 32 for 500 epochs using the Adam optimizer. To enhance convergence, we used a cyclic learning rate scheduler, which dynamically adjusts the learning rate during training. A latent space dimension of 2 is chosen for this problem.
@@
The LSTM consists of 2 hidden layers with 40 neurons[24]. The model is trained with a batch size of 32 for 400 epochs with a learning rate of $10^{-3}$ using Adam optimizer along with a cyclic learning rate scheduler. Data scaling is not performed at any stage of the training. The task here is to learn a parametric surrogate model capable of capturing the system state evolution via a reduced representation.
@@
where $\mathbf{u}=(u,v,w)$ is the velocity field, $b$ is the advection parameter and $\nu$ is the diffusion parameter. Simulation data for the 3D Viscous Burgers’ equation is obtained from APEBench dataset [34]. The advection and diffusion parameters are set to a value of -1.5 and 1.5 respectively, emulating a decaying dynamics. We generate the simulation data for 50 different trajectories, each with a different initial condition on a grid of size $32^{3}$ containing 101 temporal snapshots in each trajectory. For this dataset, we use the first 50 timesteps for training and the rest for testing purposes.
@@ 4.1 Dimensionality Reduction Performance
To highlight the state-of-the-art performance of the proposed framework, we also perform a one-off comparison of its generalization ability on the complex 3D HIT case against the $\beta$-VAE baseline [18], as detailed in Appendix C. Our framework consistently achieves superior reconstruction performance across all grid resolutions and latent space ranks. Note that the improvement in performance is obtained at little to no additional computational overhead as shown in Table 7 in terms of the total number of training parameters and the training time per epoch which be seen in Figure 13 provided in Appendix B.
@@ 4.4 Noise Robustness
Recent studies [35] in the deep learning community show that the sharpness of the minima, which describes the sensitivity of model loss with respect to perturbations in the model parameters, is a promising quantity that correlates with the generalization performance of deep networks. Let $\mathcal{D}_{\textrm{train}}=\left\{(s_{1},g_{1}),\ldots,(s_{n},g_{n})\right\}$ be the training data, that is, the set of features $s$ and target $g$ pairs, and $\ell_{i}(\mathbf{\Theta})$ be the loss of an NN model parametrized by weights $\mathbf{\Theta}$ and evaluated at the $i$th training sample point $\left(s_{i},g_{i}\right)$. Afterwards, the sharpness on a set of points $\mathcal{D}\subseteq\mathcal{D}_{\text{train}}$ can be defined as [36]: $\mathcal{S}(\mathbf{\Theta};\mathcal{D})\triangleq\max_{\|\delta\|_{2}\leq\rho}\frac{1}{|\mathcal{D}|}\sum_{i:\left(s_{i},g_{i}\right)\in\mathcal{D}}(\ell_{i}(\mathbf{\Theta}+\delta)-\ell_{i}(\mathbf{\Theta}))$ where $\delta$ is the perturbation introduced on the trainable parameters $\mathbf{\Theta}$ and $\rho$ refers to the perturbation radius. For example, Table 4 shows the sharpness and reconstruction error for a particular training instance of 1D KS data with grid points $1024$ and rank 60 and Table 5 shows the sharpness and reconstruction error for a particular training instance of 3D HIT data with grid points $32^{3}$ and rank 5. The perturbation radius used here is 0.1, which is the largest value that can be applied without causing instability in the reconstructed outputs. Our proposed framework has 1000 times less sharpness compared to the AE and simple hybrid approach in the 3D HIT case. Since the sharpness of the minima is related to the resilience of the model in the presence of noisy data, we evaluate the reconstruction performance of AE, a learnable weighted simple hybrid framework under noisy testing data. Models trained on noise-free data are tested on data that contain random normal noise with zero mean and standard deviation equal to $10\%$, $20\%$ and $30\%$ of the maximum velocity magnitude superimposed over the flow field.

### v5

@@ 1 Introduction
Computational fluid dynamics involves solving large dynamical systems with millions of degrees of freedom, resulting in significant computational overhead. In order to alleviate this problem, reduced-order modeling [deane1991low, LUCIA200451] is widely used, which uses a smaller number of modes to provide an approximate solution at a lower computational expense. A crucial step in reduced-order modeling is the projection of high-dimensional system states to a reduced latent space [taira2017modal]. The quality of projection can determine the overhead error for reduced-order modeling [rowley2005model]. Linear dimensionality reduction techniques such as Proper Orthogonal Decomposition (POD) [sirovich1987turbulence, weller2010numerical, buoso2022stabilized, mezic2005spectral] are often used to create efficient representations of large-scale systems by projecting the solution manifold onto the space spanned by a set of linear orthonormal basis.
@@
Advances in deep learning techniques, such as deep autoencoders (AE) [hinton2006reducing], capture intrinsic non-linear features for better compression, outperforming POD in low ranks, thus overcoming the well-known Kolmogorov barrier [peherstorfer2022breaking, ahmed2020breaking]. Kolmogorov barrier is defined as a slow decay of the Kolmogorov $n$-width [cohen2015kolmogorovwidthsholomorphicmappings, MADAY2002289], given by $d_{n}(\mathcal{M})=\inf_{\begin{subarray}{c}U_{n}\subset\mathcal{V},\dim(U_{n})=n\end{subarray}}\sup_{x(\cdot;t,\mu)\in\mathcal{M}}\inf_{\hat{x}\in U_{n}}\|x(\cdot;t,\mu)-\hat{x}\|$, where $\mathcal{M}$ is the solution manifold, $\mathcal{V}$ is the ambient Hilbert space, $U_{n}$ is any $n$-dimensional subspace of $\mathcal{V}$, $x(\cdot;t,\mu)$ is a solution to a parameterized PDE at time $t$ and parameter $\mu$, $\hat{x}$ is its projection onto $U_{n}$, and $\|\cdot\|$ denotes the norm in $\mathcal{V}$. The $n$-width quantifies the minimum worst-case projection error achievable by any $n$-dimensional linear subspace. This slow decay limits the best achievable error when using linear projection-based model reduction. Milano and Koumoutsakos [milano2002neural] highlighted one of the first works to utilize a fully connected autoencoder to reconstruct the flow field, offering better performance compared to POD. Further studies have reported the usage of convolutional neural networks on 2D or 3D flow fields [pawar2019deep, zhang2023nonlinear, raj2023comparison]. Such methods have been adopted in the fluid dynamics community to obtain a nonlinear model order reduction [lee2020model, solera2024beta], but they do not provide projection error convergence as the rank of the latent space increases. Recently, there have also been studies on using the hybrid approach [dar2023artificial, barnett2023neural] combining POD with deep learning, by passing the latent space produced by POD to a neural network to find the corrections required to enhance reconstruction. These hybrid techniques have proven to enhance reconstruction beyond the capabilities of vanilla autoencoders. Unlike these approaches, which treat POD as a fixed preprocessing step and use neural networks only to adjust the POD output, in this work we propose a novel dimensionality reduction technique that combines traditional dimensionality reduction technique, i.e., POD, with deep learning techniques in a weighted manner at the encoder and decoder stage, where such weights of hybridization are also learned from data, to achieve a more effective dimensionality reduction. We also compare this approach with a straightforward hybrid technique, where a direct sum of POD and AE is utilized to construct the autoencoder, demonstrating the need of using learnable weighting parameters between POD and AE. Interestingly, our proposed approach obtains flat minima as opposed to other approaches, which contributes to the improved generalization and noise robustness.
@@
Building on our proposed framework, we also demonstrate the application of our framework to surrogate modeling tasks, particularly in the context of PDE system evolution. We explore two distinct applications of our framework. First, we integrate our framework with the Koopman operator [mezic2021koopman, mezic2005spectral, arbabi2017ergodic], where our framework facilitates the mapping of Koopman embedding back to physical space, while learning a linear forward model to evolve the system states in a reduced space. Second, we leverage long-short-term memory (LSTM) for latent state evolution [hochreiter1997long]. LSTMs are usually trained on the trajectories of reduced states, obtained via POD [wang2020recurrent, mohan2019compressed] or a non-linear autoencoder [cae_lstm_rom_romit, gonzalez2018deep] to serve as a surrogate for the temporal evolution of latent states. The computational costs can be prohibitive when using LSTMs on full states, which can typically be in the range of millions for complex scenarios in fluid dynamics. By integrating our improved dimensionality reduction technique with a forward model for the latent states, we show that the overall framework leads to a more accurate prediction of the system dynamics and provide insights into the individual error contributions from the dimensionality reduction and time series model. This underscores the fact that the quality of the reduced representation is the primary bottleneck in achieving highly accurate predictions, as suboptimal dimensionality reduction inherently limit the effectiveness of any downstream models.
@@ 2.1 Dimensionality Reduction
Given the training dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{M}$, we trained the autoencoder by minimizing the mean-squared error (MSE) $\min_{\mathbf{\Theta},\mathbf{a},\mathbf{b}}\frac{1}{M}\sum_{i=1}^{M}\|\mathbf{x}_{i}-\hat{\mathbf{x}_{i}}\|^{2}$, where $\|\cdot\|$ is Frobenius norm, and $\mathbf{\Theta}$ refers to the set of the trainable parameters of neural network encoder $\phi_{\textrm{NN}}$ and decoder $\psi_{\textrm{NN}}$. $\mathbf{\Theta}$ is initialized using standard Kaiming initialization [he2015delving]. Motivated by Wang et al. [Wang2024PirateNetsPD], we choose to initialize $\mathbf{a}$ and $\mathbf{b}$ with zeros, leading to the proposed framework being equivalent to the classical POD at the beginning of neural network training. Thus, the model starts from the optimal linear encoder and becomes progressively nonlinear as the training proceeds. We choose Adam optimizer with learning rate of $10^{-4}$ for $\mathbf{\Theta}$ and $10^{-5}$ for $\mathbf{a}$ and $\mathbf{b}$.
To emphasize the role of learnable weight $\mathbf{a}$ and $\mathbf{b}$, we also implement a straightforward hybrid approach [kosut2021quantum], which simply adds the latent states from POD and NN. Similarly, the state of the reconstructed system is the sum of the output of the POD decoder and the NN decoder. We name this hybrid approach as simple hybrid autoencoder. The difference of two approaches is further illustrated in Figure 1. As we shall see in the following sections, this comparison underscores the importance of using a weighted blend of the encoded and decoded spaces, as without it, the improvement over POD is at most incremental.
@@ 2.2 Koopman Forecasting
Koopman theory [koopman1931hamiltonian] states that any nonlinear dynamical system can be linearized by lifting into a (possiblly infinite-dimensional) space of observable functions. Hence, given a continuous dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$, for any observable $\xi$, one defines the Koopman operator $\mathcal{K}$ as:
@@
To overcome these challenges, Lange et al. [lange2021fourier] proposed Koopman forecasting,
@@ 2.3 Surrogate Modeling for Time-Dependent PDEs
Our surrogate modeling strategy involves the usage of the aforementioned techniques to obtain a latent representation and LSTM to determine its evolution over time. Training is divided into two stages. The dimensionality reduction framework is first trained to obtain a low-dimensional embedding of the system states, which subsequently serves as the training data for the time series prediction model to evolve the system in time[cae_lstm_rom_romit, Halder_2024]. The advantage of training them separately is that it allows us to clearly demarcate the contribution of each stage to the overall surrogate model, which is particularly important as the crux of the current study is the introduction of a novel framework for dimensionality reduction.
@@ 3.1.2 Homogeneous isotropic turbulence (HIT)
Direct numerical simulation data of forced homogeneous isotropic turbulence is obtained from the Johns Hopkins turbulence database [li2008public]. The data set is generated by solving the forced Navier-Stokes equation on a periodic cubic box using the pseudospectral method. We interpolate the velocity field, $u(x,t)=(u_{x},u_{y},u_{z})\in\mathbb{R}^{3}$, from the original dataset of resolution $1024^{3}$ to grid sizes of $16^{3}$, $32^{3}$ and $64^{3}$ and extract the data from timestep 1 to 2048 with a stride of 16, resulting around 128 snapshots.
We choose a deep convolutional autoencoder [fukami2020convolutional] as the NN part of the proposed framework. The details of the architecture pertaining to the NN based encoder and decoder and the training hyper parameters are shown in Table 1. Without shuffling, we systematically sample the dataset by selecting every alternate snapshot to curate the training and testing data, each comprising 64 snapshots.
@@ 3.2.2 Flow Over Cylinder
We utilize the time evolution data of vorticity in a two-dimensional flow over a cylinder at $Re\approx 100$ [lange2021fourier]. The initial 50 temporal snapshots are used in training the Koopman-Decoder model, and the remaining 100 are kept out for testing. The details of the non-linear part of the decoder and training parameters are provided in Table 3. The Koopman model is parametrized by two frequencies, essentially evolving the dynamics in a latent space of rank 4.
@@ 3.3 Surrogate Modeling Datasets
A fully connected network using the same architecture as outlined in Section 3.1.1 is used for the NN based encoder and decoder. The dimensionality reduction framework was trained with a learning rate of $1\times 10^{-3}$, batch size of 32 for 500 epochs using the Adam optimizer. To enhance convergence, we used a cyclic learning rate scheduler, which dynamically adjusts the learning rate during training. A latent space dimension of 2 is chosen for this problem.
@@
The LSTM consists of 2 hidden layers with 40 neurons [cae_lstm_rom_romit]. The model is trained with a batch size of 32 for 400 epochs with a learning rate of $10^{-3}$ using Adam optimizer along with a cyclic learning rate scheduler. Data scaling is not performed at any stage of the training. The task here is to learn a parametric surrogate model capable of capturing the system state evolution via a reduced representation.
@@
where $\mathbf{u}=(u,v,w)$ is the velocity field, $b$ is the advection parameter and $\nu$ is the diffusion parameter. Simulation data for the 3D Viscous Burgers’ equation is obtained from APEBench dataset [koehler2024apebench]. The advection and diffusion parameters are set to a value of -1.5 and 1.5 respectively, emulating a decaying dynamics. We generate the simulation data for 50 different trajectories, each with a different initial condition on a grid of size $32^{3}$ containing 101 temporal snapshots in each trajectory. For this dataset, we use the first 50 timesteps for training and the rest for testing purposes.
@@ 4.1 Dimensionality Reduction Performance
To highlight the state-of-the-art performance of the proposed framework, we also perform a one-off comparison of its generalization ability on the complex 3D HIT case against the $\beta$-VAE baseline [solera2024beta], as detailed in Appendix C. Our framework consistently achieves superior reconstruction performance across all grid resolutions and latent space ranks. Note that the improvement in performance is obtained at little to no additional computational overhead as shown in Table 7 in terms of the total number of training parameters and the training time per epoch which be seen in Figure 13 provided in Appendix B.
@@ 4.4 Noise Robustness
Recent studies [foret2020sharpness] in the deep learning community show that the sharpness of the minima, which describes the sensitivity of model loss with respect to perturbations in the model parameters, is a promising quantity that correlates with the generalization performance of deep networks. Let $\mathcal{D}_{\textrm{train}}=\left\{(s_{1},g_{1}),\ldots,(s_{n},g_{n})\right\}$ be the training data, that is, the set of features $s$ and target $g$ pairs, and $\ell_{i}(\mathbf{\Theta})$ be the loss of an NN model parametrized by weights $\mathbf{\Theta}$ and evaluated at the $i$th training sample point $\left(s_{i},g_{i}\right)$. Afterwards, the sharpness on a set of points $\mathcal{D}\subseteq\mathcal{D}_{\text{train}}$ can be defined as [andriushchenko2022towards]: $\mathcal{S}(\mathbf{\Theta};\mathcal{D})\triangleq\max_{\|\delta\|_{2}\leq\rho}\frac{1}{|\mathcal{D}|}\sum_{i:\left(s_{i},g_{i}\right)\in\mathcal{D}}(\ell_{i}(\mathbf{\Theta}+\delta)-\ell_{i}(\mathbf{\Theta}))$ where $\delta$ is the perturbation introduced on the trainable parameters $\mathbf{\Theta}$ and $\rho$ refers to the perturbation radius. For example, Table 4 shows the sharpness and reconstruction error for a particular training instance of 1D KS data with grid points $1024$ and rank 60 and Table 5 shows the sharpness and reconstruction error for a particular training instance of 3D HIT data with grid points $32^{3}$ and rank 5. The perturbation radius used here is 0.1, which is the largest value that can be applied without causing instability in the reconstructed outputs. Our proposed framework has 1000 times less sharpness compared to the AE and simple hybrid approach in the 3D HIT case. Since the sharpness of the minima is related to the resilience of the model in the presence of noisy data, we evaluate the reconstruction performance of AE, a learnable weighted simple hybrid framework under noisy testing data. Models trained on noise-free data are tested on data that contain random normal noise with zero mean and standard deviation equal to $10\%$, $20\%$ and $30\%$ of the maximum velocity magnitude superimposed over the flow field.